# CrowdGenUI: Enhancing LLM-Based UI Widget Generation with a Crowdsourced Preference Library

### v1

@@ 7.4.3. Reflection on Research Questions
For RQ2, the results demonstrate that incorporating the crowdsourced library enhances the ability of the LLM to reason widgets that align with user preferences regarding predictability, efficiency, and explorability. In terms of predictability, users’ preferences captured from the crowdsourcing study lean towards preset buttons for predefined value selection, color wheels, and color pickers for color adjustments, and click on image for position changes (recall Section 6.3.2). These preferences have directly influenced the widgets reasoned for the user evaluation tasks, where widgets reasoned with the library adhered more closely to these favored widget types. In contrast, when the library is not used, the LLM reasons sliders for most tasks regardless of the task categories. For efficiency, similar patterns emerge—preset buttons, color wheels, color pickers, and click on image are repeatedly preferred for their ease and speed in achieving task goals. The LLM, when equipped with the crowdsourced library, is able to reason toward these efficient widgets, whereas, without the library, the reasoning leans more heavily on sliders and preset buttons for most tasks without too much differentiation for task categories. Regarding explorability, the user preference trends show that sliders are favored for continuous value adjustments, color wheels are preferred for color changes, and click on image is liked for position adjustments. Without the library, however, the LLM tends to reason sliders to allow exploration for continuous value adjustments, color pickers for color changes, and sliders and text fields for position adjustments. The incorporation of the crowdsourced library enables the LLM to capture the need for exploration in tasks like color change and position adjustment by effectively aligning the widget choices with user preferences.
@@
For RQ3, we observe that withlib30, containing the most crowdsourced data, emerges as the most preferred library across participants, particularly for predictability and explorability, while performing almost equally with withlib25 in terms of efficiency. This pattern, supported by the distribution of user preferences across all tasks, as shown in Figure 12, highlights withlib30 as the most reliable for all three aspects due to the richness of its crowdsourced data. This consistency of user preferences underscores the value of extensive, diverse data in reasoning well-aligned widget options for users. On the other hand, withoutlib shows more significant variability in user preferences, particularly in predictability and efficiency, suggesting that when no library is used, LLM reasoning lacks the consistency that is otherwise afforded by crowdsourced data. While withlib10 and withlib25 are moderately favored for predictability and efficiency, they do not perform as well in explorability as withlib30, further emphasizing the benefit of larger datasets in supporting widget reasoning for tasks requiring exploration. However, examining each task individually reveals significant variation in preferred libraries for different preference aspects. This indicates that while larger libraries lead to consistent preference-aligned widget reasoning across multiple tasks in general, for individual tasks, the utility of LLM-reasoned widgets for specific preference aspects may need additional evaluation.
Figure 12. Distribution of user preference selection counts of three preference aspects across all tasks.
@@ 8. Discussion
Based on the analysis of user evaluation results, this section discusses the implications of these results from three perspectives: the balance between task-generic and task-specific LLM widget generation, the alignment with user preferences for user-centered UI design, and the use of user preference data to create UIs that adapt to personalized use cases and evolve over time.
8.1. Generalizability of User Preference Data to Support Various Tasks
The user evaluation findings suggest that crowdsourced UI widget preference libraries provide an effective foundation to improve LLM reasoning across tasks beyond the original scope of crowdsourcing. From the tasks tested in the user evaluation, we have shown the task generalizability not only in the image editing domain but also in the design domain (recall tasks 5 and 6 in Table 2). This generalization ability can reduce the need for task-specific widget customization using automatic LLM widget generation, streamlining the UI design process. To further enhance this generalizability, the scope of crowdsourced tasks can be expanded by incorporating a more comprehensive array of task categories beyond continuous value adjustment, discrete value selection, and color and position changes. In addition, we can deepen the exploration within specific task categories to involve more complex aspects of data manipulation. For instance, our study emphasized tasks that required exploration and precision within the three task categories, and this can be extended to tasks on content editing or value adjustment that emphasize simplicity, speed, or flexibility. Also, while we focused on single-value adjustment, this setup can be extended to multi-dimensional data manipulation or compound parameter adjustments. By carefully curating the tasks for crowdsourcing, we envision that LLMs can be equipped with crowdsourced insights to reason UI widgets in both task-generic and task-specific contexts, adapting to diverse user needs and task scenarios.
Moreover, combining task-generic and task-specific UI widget reasoning could provide a solution for cases where generalized user preferences alone are insufficient. In our user evaluation, preference aspects of a few tasks showed that user preferences were better supported without the library (recall Sections 7.4.2 and 7.4.3). These findings highlight that crowdsourced preferences do not universally apply to all tasks, and some require more refined, context-sensitive reasoning. To address this, an adaptive approach that integrates task-specific customization with the general crowdsourced library can potentially better accommodate unique user interaction demands to ensure the LLM delivers more tailored and effective UI widget suggestions across a wide range of tasks. For example, for the efficiency aspect of image_adjust_tint, users showed a preference for preset buttons generated by withoutlib over the color wheel and color picker reasoned with the library. Incorporating this insight into the LLM reasoning process by assigning higher weights to preset buttons for tasks involving image tint adjustments and related operations can lead to more preference-aligned UI generation than directly using the library. This approach leverages the flexibility of LLM-based widget generation by taking updated user insights to adapt to both generic and specific task requirements.
8.2. Towards User-Centered UI Widget Selection and Design
The crowdsourcing results reveal significant variation in user preferences for UI widgets based on different data manipulation needs, emphasizing the importance of a user-centered approach to UI widget selection. This approach contrasts with the traditional designer-dominant method, where UI widgets are often pre-selected by software designers. For instance, tools like Photoshop provide a slider and a text field for adjusting image lightness, saturation, and hue, which offer a one-size-fits-all solution. However, our crowdsourced data shows that users prefer distinct widgets for various data manipulation requirements, such as preset buttons for efficient adjustment of lightness and saturation and a color wheel to explore multiple hue options (recall Section 6.3.2). This divergence reveals a gap between user-preferred and designer-selected UI widgets. This gap highlights that static, predefined UI widgets in most software may not always meet the diverse needs of users. LLM-supported UI generation offers a promising solution to bridge this gap by generating dynamic widgets. With the crowdsourced preference library, we have extended the existing LLMs’ ability to reuse user-preferred widgets based on user preference insights and, therefore, generate preference-aligned widgets for a broad range of tasks. This framework can be used to enable the creation of user-centered interfaces that align with the preferences collected from users to augment the traditional UI widget selection method that heavily relies on designers.
Furthermore, the user-centered UI design approach can be expanded to accommodate a broader spectrum of preference aspects and cater to more diverse user groups. Our study focuses on three aspects of user preferences, demonstrating the prototype’s effectiveness in guiding LLM reasoning, but future research could incorporate additional dimensions of UI design. For example, prior work on UI design principles (McKay, 2013; Lidwell et al., 2010; Chen et al., 2021), which covers a number of guidelines, can be used to inform crowdsourcing efforts. Among them, accessibility is a particularly important dimension that can be addressed by collecting user preferences from individuals with disabilities. Crowdsourced data from these users could inform the design of accessible UI widgets, ensuring that interfaces meet their specific needs. This method would not only aid LLMs in generating accessible widgets that benefit users but also provide valuable insights for software designers. For example, LLMs can be equipped with crowdsourced data to suggest widget sizes optimized for low-vision users or alternative interaction techniques for those with motor impairments. They can also be prompted to offer reasonings for their suggestions, which could serve as references for design rationale to assist designers. Thus, the LLM-reasoned widgets can potentially augment the ideas proposed in prior research, such as SUPPLE (Gajos and Weld, 2004) and SUPPLE++ (Gajos et al., 2007), to develop dynamic UIs that accommodate motor and vision capabilities.
8.3. Opportunities for Adaptive UIs
The comparison between different library sizes reveals that larger libraries tend to produce more consistent preference-aligned widget reasoning across various tasks; meanwhile, it highlights the potential of smaller libraries for cost-effective and adaptive UI design. Our study shows that even with a library of just 10 user responses, i.e., withlib10, key insights into user preferences can be effectively captured. In some tasks and aspects, such as the predictability of adjusting image exposure and tint, the efficiency of changing colors to spring tones and positioning a logo, and the explorability of aligning text and adjusting tint, smaller libraries performed as well or even better than larger ones. Small-scale libraries are quick to collect and easy to evolve, providing efficient UI design suggestions for designers to reference in the design process. Even at the extreme, i.e., using no library at all, LLMs can reason UI widgets that, in some instances, outperformed those based on larger libraries. This demonstrates that LLMs can aid in UI widget reasoning and selection without requiring extensive human input, which offers designers a low-cost way to access user preferences during UI design. Such adaptability of small libraries can significantly reduce the time and resources required to gather large datasets while still yielding meaningful user-centered insights.
The adaptability enabled by small-scale libraries for UI design can be extended to deeper UI personalization, which reflects individual preferences and supports long-term UI evolution. By collecting a few samples from individual users, these libraries can be tailored to represent personalized preferences for UI widget generation. Such personalization can allow users to build workflows that align with their specific needs. Moreover, personalized libraries have the potential to evolve alongside users as their skills and expertise grow. For example, the LLM could initially generate simple, beginner-friendly widgets and, over time, adapt to offer more advanced, feature-rich widgets as the user becomes more proficient by using the user’s updated personal library it uses for widget reasoning. This dynamic evolution enables users to progressively engage with more complex tools to support an adaptive and long-term learning curve. This kind of UI evolution not only streamlines immediate tasks but also fosters continuous improvement in tool use, helping users to gradually master more sophisticated functions without overwhelming them at the start.
@@ 9. Limitations and Future Work
The first limitation lies in the range of preferences covered during crowdsourcing. Our study focused on three key aspects, but there are other significant UI design principles, such as accessibility, cognitive load, and adaptability (Lidwell et al., 2010), that were not explicitly addressed here. However, our framework is highly extendable and can be enhanced by incorporating a broader set of preference dimensions with customizable aspects to account for user preferences from different perspectives.
Next, we mainly considered software-based UI widget generation in this work. While these widgets suited our tasks, other potential options, like gesture-based or tangible widgets, were not included. This may limit our understanding of how different widgets, like those supporting spatial interactions in XR (Aponte et al., 2024) or physical interactions in tangible interfaces (Campos Zamora et al., 2024), could meet specific task needs or user preferences. Expanding the widget types in future research could offer comprehensive insights into user preference across a broader spectrum of tasks and interaction styles.
Lastly, our system implementation is currently focused on UI generation within the image editing domain. In the future, we aim to collect data from multiple domains and various tasks to build a more extensive user preference library. This would allow us to evaluate how a cross-domain library can improve widget generation, both for in-domain tasks and for tasks outside the collected domains. Such a system would also introduce challenges, such as how to effectively manage large datasets and ensure the LLM fully leverages all the available knowledge.

### v2

@@ 6.5.2. Preference Alignment of Generated UI Enhanced by Crowdsourced Library [RQ2]
The results demonstrate that incorporating the crowdsourced library enhances the LLM’s ability to reason widgets that align with user preferences. For predictability, users’ preferences captured from the crowdsourcing lean towards preset buttons for predefined value selection, color wheels, and color pickers for color adjustments, and click on image for position changes (recall Section 5.4). These preferences have directly influenced the widgets reasoned for the user evaluation tasks, where widgets reasoned with the library adhered closely to these favored widget types. In contrast, when the library is not used, the LLM reasons sliders for most tasks regardless of the task categories. For efficiency, similar patterns emerge—preset buttons, color wheels, color pickers, and click on image are repeatedly preferred for their ease and speed in achieving task goals. The LLM, when equipped with the crowdsourced library, can reason toward these efficient widgets, whereas, without the library, the reasoning leans more heavily on sliders and preset buttons for most tasks without too much differentiation for task categories. Regarding explorability, the user preference trends show that sliders are favored for continuous value adjustments, color wheels are preferred for color changes, and click on image is liked for position adjustments. Without the library, however, the LLM tends to reason sliders to allow exploration for continuous value adjustments, color pickers for color changes, and sliders and text fields for position adjustments. The incorporation of the crowdsourced library enables the LLM to capture the need for exploration in tasks like color and position adjustment by aligning the widget choices with user-preferred ones.
Takeaway: Integrating the crowdsourced library enables the LLM to reason widgets that more accurately align with user preferences for predictability, efficiency, and explorability, while its absence leads to more generic and less tailored reasoning.
6.5.3. Effect of Library Sizes on UI Generation [RQ3]
@@
Both the total selection counts and their distributions, shown in Figures 9(a) and 9(b), highlight withlib30 as the most reliable for all three aspects due to the richness of its crowdsourced data. This consistency of user preferences underscores the value of extensive, diverse data in reasoning user-aligned widgets. On the other hand, withoutlib shows significant variability in user preferences, suggesting that when no library is used, LLM reasoning lacks the consistency that is otherwise afforded by crowdsourced data. However, examining each task individually reveals some variation in preferred libraries for different preference aspects. This indicates that while larger libraries lead to consistent preference-aligned widget reasoning in general, tasks like image_adjust_temperature and image_adjust_tint show that withoutlib is almost equally favored as withlib30 regarding certain aspects and may need additional evaluation.
Takeaway: Extensive crowdsourced data makes it the most reliable and consistent for widget reasoning across tasks, while libraries with fewer data lead to greater variability in user preferences.
@@ 7. Discussion
In this section, we reflect on our findings and discuss their implications for crowdsourcing-enhanced human-centered UI generation.
7.1. Improving Task Generalizability with Curated Crowdsourcing Setup
Importance of task curation to collect helpful user data that can boost LLM-based UI generation: The user evaluation findings indicate that the crowdsourced user preference library improved the LLM’s reasoning capabilities, extending the library’s effectiveness beyond its initial intended task scope and domain. Such generalizability highlights the user data’s potential to serve as a foundational element to guide the creation of versatile, human-centered UIs across diverse domains through careful task curation. This can contribute to a significant reduction in the necessity for task-specific UI customizations. For instance, if we collect user-preferred UIs to adjust image tones, this insight can be helpful in indicating user preferences for a range of similar tasks involving continuous color adjustment. In our implementation, the effectiveness of incorporating tasks belonging to the three categories involving diverse data manipulation requirements (recall Section 5.3.2) has been demonstrated through our evaluation. This implies that through well-curated task design, UI designers can collaborate with LLMs that are equipped with user data to prototype designs to meet users’ expectations with reduced additional user testing.
Extending the breadth and depth of task categories: To further enhance the user data’s generalizability, future efforts can strategically encompass a broader spectrum of task categories. In addition to this work’s focus on image parameter adjustment, other potential task categories include audio, video, data visualization editing, and others introduced in Section 2. These domains often involve distinct interaction styles. For instance, video editing might require precise temporal controls like scrubbers and frame-level sliders, while data visualizations might prefer touch-based navigation on tablets versus hover-driven interactions on desktops. Because each domain brings unique data manipulation, their UI designs may not be easily generalizable across domains. As such, multi-domain user data collection is essential to ensure UI recommendations remain relevant and effective for diverse task contexts.
On the other hand, future research can also aim for a deeper exploration of specific task categories, focusing on more complex and intricate forms of data manipulation. For instance, for image editing, we can extend the scope from single-value to multi-dimensional or compound parameter manipulations. This is a common need as image editing often consists of a combination of multiple parameters, e.g., if a user wants to correct a low-exposure image, they may increase both image saturation and brightness. Thus, a dataset containing user data on complex or open-ended tasks may help inform the LLM-based UI generation framework to create a widget that combines multi-parameter adjustment or to generate widget combinations. Investigating extended task complexities can help uncover the dependence of specific parameters and the relevance of high-level task requirements, thus refining the granularity of crowdsourced data to bolster the applicability of LLM-generated UIs in addressing complicated task contexts.
7.2. Designing User-Focused UIs with Insights Collected from Target Audience
Tailoring UIs to target users favored over traditional fixed UI designs: Our crowdsourcing results highlight variations in user preferences for UI widgets, underscoring the need for a user-centered approach to widget selection. This contrasts with traditional ways of UI design by most software and applications, i.e., UIs are fixed and consist of designers’ pre-selected UI elements. For instance, Photoshop offers a slider and text field for adjusting image lightness, saturation, and hue. However, our crowdsourced data reveals various user-preferred widgets that go beyond these given options. This divergence exposes a gap that static, designer-selected widgets sometimes fail to meet diverse user needs. To this end, our framework utilizes LLM-based dynamic UI design, which generates UIs aligned with user preferences and task contexts. This framework facilitates the creation of user-centered interfaces with direct end-user input versus the traditional designer-driven methods.
Collecting specific, targeted user data to guide tailored UI design: Our user-centered framework can further evolve to encompass a broader range of preference dimensions and serve more diverse user populations. While this work uses three preference aspects to demonstrate the framework’s efficacy, future work can explore additional dimensions by using established UI design principles (McKay, 2013; Lidwell et al., 2010; Chen et al., 2021; Shneiderman, 2000; Norman, 1983). Among them, we take accessibility as an example to show how other dimensions can be incorporated. Accessible UI design is a crucial dimension and is addressable through crowdsourcing preferences from users with disabilities. The collected data would enable LLMs to reason accessibility considerations directly based on the target user’s data. For instance, users with visual impairments might prefer high-contrast color schemes, screen reader-friendly labels, or voice-based navigation. Those with motor impairments may favor interfaces with larger clickable targets, simplified gesture controls, or customizable input methods based on varied motor abilities (Gajos and Weld, 2004; Gajos et al., 2007). These accessibility preferences, once incorporated into the user library, can be used to guide LLM reasoning during UI generation. In turn, the LLM can recommend design adaptations tailored to specific user groups, such as suggesting larger widget sizes or reducing the need for precise cursor movements, and provide explanations grounded in the collected data. This not only supports more inclusive UI outcomes for end users but also makes the rationale behind design recommendations transparent for designers’ reference.
7.3. Building Adaptive UIs with Scalable User Data
Multi-scale user data to guide LLM reasoning on UI designs: Based on the comparisons of libraries of varying sizes, our results reveal that larger libraries generally yield consistent preference-aligned widget reasoning across diverse tasks. On the other hand, smaller libraries stand out for their cost-effectiveness and adaptability to guide UI design. Our study demonstrates that a library with as few as 10 user responses can effectively capture insights into user preferences. In specific tasks, such as adjusting image exposure, shifting colors to spring tones, and positioning logos, small-scale libraries perform comparably with their larger counterparts for certain preference aspects. These libraries are quick to compile, offering users and designers fast UI design recommendations without the need for extensive data collection and user testing. This highlights how LLMs can support UI design with cost-efficient human input for user-centered design.
Low-cost, small-scale user data for UI personalization and evolution: The adaptability of small-scale libraries can enable deeper personalization and long-term UI evolution. By incorporating a few user-specific samples, the small data can be used to reflect individual preferences and tailor workflows that align with unique needs. Such UI personalization and evolution can be especially beneficial for beginner users who want to use professional, feature-rich software without feeling overwhelmed at the start. As users’ skills and expertise develop, their personal data can evolve and dynamically adjust the UIs to be more advanced over time. For instance, we can use our framework to initially generate simple, beginner-friendly UIs, then progressively adapt to offer sophisticated UIs with advanced tools as the user’s proficiency grows, guided by an updated personal library covering user traces. This evolution monitors user growth and fosters a gradual and personalized learning curve, allowing users to engage with increasingly complex tools at their own pace. With a similar approach, creators can also use UI personalization and evolution to tailor UIs to individual workflows. Although existing software allows users to show/hide tools for UI personalization (Adobe, 2024c; Blender, 2024), this often does not allow transformative UI redesign like regrouping tools into a user-created panel. Instead, using AI to generate UIs can programmatically create tools with personalized, evolving functionality that is tailored to creators’ own needs. This dynamic approach can be beneficial to enhance productivity as users are not required to learn and adapt to given UIs, but instead, UIs are tailored to their workflows.
7.4. Exploring UI Design in Human-Centered Agentic AI
Indispensable role of UIs in agentic AI: As agentic AI increasingly automates complex tasks, e.g., building agents with Magentic-One (Fourney et al., 2024) to write and execute code, we envision agentic AI systems to help delegate tasks that can significantly foster human productivity. In the meantime, alongside the pursuit of fully autonomous agentic AI systems, designing human-AI collaboration systems remains a critical problem in balancing efficiency with human agency. While agentic systems can autonomously execute actions on behalf of humans, their effectiveness hinges on aligning with user expectations and contextual needs. This challenge is particularly pronounced in domains requiring creativity and exploration. Taking image editing as an example, a task involving image parameter adjustment can yield numerous solutions that different creators come up with. Thus, agentic AI that automates the parameter adjustment can easily fall short of satisfying such subjective and aesthetic needs and lead to numerous iterations of prompt refinement. Instead, UIs with direct controls that bridge the communication between humans and agentic AI become essential. Thus, we argue that the user interface remains indispensable in agentic AI systems where meaningful human engagement is desired.
Dataset injection to LLMs as a scalable solution for UI generation to support human-centered agentic AI: To allow interface interaction in agentic AI, LLM-generated UIs present a dynamic solution well-suited to the automation and adaptability required by LLM-based AI agents. However, using LLMs trained on generic data alone risks homogenizing design choices to generic solutions unless explicitly guided by target user-driven insights, as shown in this work. This highlights the need for UI generation frameworks to prioritize user needs. Among existing LLM post-training techniques for customizing LLMs to specific tasks and user preferences (Kumar et al., 2025; Tie et al., 2025), our framework contributes an approach of embedding a crowdsourced user dataset to guide LLMs to produce user-centric outputs through structured reasoning enhanced by this dataset. More interestingly, this framework can further benefit from agentic AI systems to deliver UI designs leveraging agent collaboration. Imagine that we can have multiple LLMs working simultaneously to generate UI designs, each equipped with user data covering different user profiles and design principles. A multi-agent system can be built to analyze these LLMs’ outputs, coordinate conflicts, and produce UI design suggestions with multiple perspectives considered. This agentic collaboration can enhance coverage and diversity in UI design, as well as support dynamic dialogue between agents to enable emergent design creativity.
@@ 8. Limitations and Future Work
New interaction modalities beyond traditional GUIs: Although our work primarily focused on 2D UI widget generation, an emerging area is applying the framework to novel interaction modalities beyond 2D GUIs for preference alignment. For instance, as voice-controlled interfaces, AR/VR environments, and wearable device UIs become increasingly popular, it would be interesting to collect user preferences for those contexts. Future work can crowdsource user data regarding the voice command structure for a given function based on what most children find intuitive, gestures, or holographic controls that most senior people feel natural in AR, and gesture-based shortcuts or low-interruption interfaces for smartwatches suitable for most people in office settings. This direction can extend the crowdsourcing-informed interaction design as UIs diversify with technology advancements and user contexts.
Whole-interface generation and layout: This work focuses on individual UI widget generation, and future research can expand to the complete interface layout or workflow generation by investigating how a crowdsourced dataset scales to decisions about arranging multiple widgets, navigation structures, or multi-step task flows. For example, given a complex task (e.g., “design a photo editing toolbar”), what interface layouts do users prefer, and how can an LLM assemble those using collected user insights? Crowdsourcing could be used to gather user-preferred layouts or sequences in addition to widget types. This can extend LLM-based UI generation to entire UIs that are human-centered at a macro level and also opens questions about representing and learning from more complex user data, e.g., sequences of interactions and grouping of controls, and ensuring coherence in generated designs.
Integration with design guidelines and experts: Future work can combine the current crowdsourced general user data with established HCI design principles or input from professional UI designers. Research might explore hybrid systems where the LLM is guided by both the crowd-derived library and rule-based heuristics or where crowd preferences are augmented with expert critiques. This addresses whether certain nuanced design decisions (e.g., accessibility considerations or aesthetic consistency) might require more than just crowd preferences. By involving experts in the loop, e.g., crowdsourcing from designers or using expert-verified crowd data, the resulting UI generation could achieve a balance between general user appeal and professional best practices.
Cross-cultural user data collection: The current work crowdsourced user preferences primarily from English-speaking participants. This can lead to overlooked preferences and cultural nuances that emerge in broader global contexts. For UI design, user preferences for color schemes, layout conventions, or even interaction metaphors may differ substantially across linguistic and cultural divides. Future data collection could employ multilingual crowdsourcing platforms to enhance cross-cultural inclusiveness and adapt tasks to reflect culturally relevant use cases. This would allow LLMs to account for cross-cultural insights to generate UIs that reflect the diversity of global user populations, thus promoting equitable and context-aware UI design on a global scale.
User persona-aware UI design: Beyond user preferences, future research can incorporate broader user persona data, such as demographic information (e.g., age, gender), domain expertise (e.g., novice vs. expert users), device usage habits, cognitive styles, and other aspects, to represent essential aspects of target users. By leveraging these diverse persona attributes, the framework can be extended to process persona data and further generate UI designs that are not only preference-aligned but also contextually appropriate for different user types.
Technical advancements of using LLMs for UI generation: This work implements a system prototype of the proposed framework and opens up opportunities for technical improvements. First, using LLMs for code generation introduces limitations, as minor bugs may occur and require manual fixes. Recent research has developed LLM-based agents to fix bugs in code automatically (Yang et al., 2024; Zhang et al., 2024; Kang et al., 2025). This offers a promising solution to robustly generate UIs in future iterations. Next, we use GPT-4o as the LLM for UI generation, and future research could explore more advanced LLMs for improved reasoning capabilities, open-source LLMs for lower cost and easier replicability, or fine-tuned LLMs for specific task domains. This would also lead to an in-depth comparison of different LLMs’ performance for UI generation.